a particle is projected vertically upwards prove that it will be at 3/4of hight H at times which are in ratio 1:3

Grade:11

1 Answers

Aditya mishra
5 years ago
Given t/t'=1/3
 
 t=1sec.         t'=3sec.          t-t'=2 sec the time at which particle is at its top most point where velocity is 0
   V= u. +  at>>>>>>>>>>>>> U. = 20 m/s
 
Height of top most point is H=u²/2g
 H= 20 meter
Displacement of the particle after 1sec 
h= ut -1/2at²>>>>>>>>>>>= 15 meter 
displacement of the particle after 3 sec h' = 15 meter 
 
h = h'= 15 =¾×15= ¾H